### v1

# The Essential Role of Causality in Foundation World Models for Embodied AI

Abstract
Recent advances in foundation models, especially in large multi-modal models and conversational agents, have ignited interest in the potential of generally capable embodied agents. Such agents would require the ability to perform new tasks in many different real-world environments. However, current foundation models fail to accurately model physical interactions with the real world thus not sufficient for Embodied AI. The study of causality lends itself to the construction of veridical world models, which are crucial for accurately predicting the outcomes of possible interactions. This paper focuses on the prospects of building foundation world models for the upcoming generation of embodied agents and presents a novel viewpoint on the significance of causality within these. We posit that integrating causal considerations is vital to facilitate meaningful physical interactions with the world. Finally, we demystify misconceptions about causality in this context and present our outlook for future research.
Machine Learning, ICML
1 Introduction
The ability to perform physically meaningful interactions in a real-world environment is essential to maximize the potential of technology in everyday life. Entities capable of conducting interactions of this kind are in our work referred to as embodied agents. These agents can vary in form, ranging from robots to humans augmented by AI in mixed reality environments, or in virtual reality where interactions bear physical importance.
Foundation models have recently accounted for the most significant advancement in generalization across a range of tasks with multi-modal input, compared to canonical deep learning models from the past decade. In particular, a single model learns to address a wide variety of tasks in a zero-shot manner. In robotics, we anticipate that foundation models for embodied agents will make the introduction of new robots into our life effortless: each new robot can immediately solve a given task using the same model without the need for a specialized model. In addition, robots can quickly adapt to a new environment by interacting with it and use the new experiences to improve without lengthy retraining or complex programming.
However, contemporary foundation models are not sufficient. Current approaches, dominated by large (vision-) language models (Achiam et al., 2023; Bubeck et al., 2023; Chen et al., 2023), are based on correlational statistics and do not explicitly capture the underlying dynamics, compositional structure or causal hierachies. The lack of a veridical world model renders them unsuitable for use in embodied AI, which demands precise or longterm action planning, efficient and safe exploration of new environments or quick adaptation to feedback and the actions of other agents.
Causality at its core aims to understand the consequences of actions, allowing for interaction planning. Philosophy and cognitive sciences sees understanding causal concepts as both the basics and sometimes the ultimate goal for humans to learn how to interact with the world (Gibson, 1978; Gopnik et al., 2007; Adams & Aizawa, 2021) and key in children’s development (Piaget, 1965; Gibson, 1988). This aligns with the concept of affordances in Embodied AI research (Kjellström et al., 2011; Koppula et al., 2013; Koppula & Saxena, 2013; Ardón et al., 2020; Ahn et al., 2022), which pertains to the types of action that can be applied and the resulting consequences. Importantly, even with the help of available real or simulated environments, the experimentation might still be too coarse-grained to deal with spurious relationships (Herd & Miles, 2019; Lavin et al., 2021). Thus, the world model for an agent, including humans, should be causal-aware (Anonymous, 2024b) and able to efficiently incorporate data from demonstration and interaction (i.e., the property referred as veridicality in this paper). As such, causality offers tools and insights that hold the key pieces to building Foundation Veridical World Models (FVWMs) (Section 2.2) that will power future embodied agents.
In this position paper, we argue that future progress in Embodied AI relies on advances in foundation models, given the readiness of various technological considerations (Section 2). For such research, causality is pivotal in developing the FVWM, which requires us to go beyond canonical causal research paradigms (Section 3). Thus, we foresee research opportunities (Section 4) advancing foundation world models in Embodied AI from a causal perspective. We extend our discussion with considerations in deploying Embodied AI (Section 5) in the end.
2 Embodied AI and Foundation Veridical World Models
Foundation Veridical World Model Daily assistance humanoid Industrial manufacturing robots Mixed-reality interactions … …
Figure 1: The foundation world model is designed to comprehend associations, counterfactuals, and interactions within the world, aiding any embodied AI agent in executing tasks across diverse environments and platforms.
For many years, research has pursued the creation of real Embodied AI. Yet, the tangible success of these technologies is predominantly seen in scenarios where environments are tailored to suit the robot for a particular task, such as assembling machines in production lines. The inability to adapt to diverse environments and multiple tasks limits their real-world applicability and the range of problems it can solve. It highlights a significant gap in fulfilling the promises of embodied AI to be a multi-task solver in general environments. We believe that recent technological advancements (Section 2.1), including hardware, foundation models, and data, have brought general Embodied AI within reach, where the remaining gap can be overcome by veridical world models (Section 2.2).
2.1 The readiness of Embodied AI research
Hardware advancement
The pace of innovation in hardware and systems, ranging from real-world robots to virtual and mixed reality devices, is advancing at an unprecedented rate, though there is still a significant journey ahead before achieving maturity for general tasks. The costs of both humanoid (e.g. Optimus, Phoenix (Sanctuary, 2024), Digit (Agility, 2024)) and quadrupedal robots (e.g. Spot (BostonDynamics, 2024), Go2 (Unitree, 2024) ) have been rapidly decreasing over the last decade, with a huge potential market size (GoldmanSachs, 2023; Tilley, 2017). For virtual mixed reality experiences, Oculus from Meta has increased in popularity and Apple’s Vision Pro is taking a commercial leap offering new interactive experiences, creating opportunities for virtual agents. In addition, new types of sensors for robots have been invented recently. Electronic skin can capture the sensation of touch (Sanderson, 2021) to enable dexterous robots (Lambeta et al., 2020; Sun et al., 2022), while perovskite retinomorphic sensors (Trujillo Herrera & Labram, 2020) mimic the human retina and react to changes in illumination rather than constant signals. These advancements offer different ways to sense the world. This enables the consideration of more challenging manipulations, such as those involving deformable objects, rather than rigid ones (Yin et al., 2021).
Meanwhile, the evolution of AI computing power has accelerated at an unprecedented rate. This surge in computing capabilities is primarily driven by innovations in hardware, such as more powerful GPUs and specialized AI processors (Dally et al., 2021; Saravanan & Kouzani, 2023), providing the computing basis for foundation models. These developments have enabled AI systems to process and analyze vast amounts of data more quickly and more effectively than ever before, thereby expanding the horizons of potential across various domains.
Paradigm shift in AI
The paradigm shift driven by foundation models in both language and vision has markedly enhanced the AI’s ability to understand and perceive the environment (Bubeck et al., 2023; Achiam et al., 2023; Yuan et al., 2021; Liu et al., 2023; Chen et al., 2023). These systems now demonstrate an increasing capability to generalize on high-level planning and decision-making tasks, often in a zero-shot manner. This has led to a shift in the research paradigm from specialized models for a single target task and dataset to foundation models generalizing across tasks and multiple datasets. The increased generalization capabilities hold the key to build foundation veridical world models required for embodied AI, opening the doors to applications that were once confined to the realms of imagination.
Data availability
Adding to the timeliness of this venture is the increasing amount of of data in the field of embodied AI, which is crucial for effective training of foundation world models. Robotics simulation environments and demonstration datasets have become more diverse, encompassing different types of tasks and scenarios, such as OpenX embodiment (Open X-Embodiment Collaboration et al., 2023), Isaac gym (Makoviychuk et al., 2021; Ma et al., 2023), CloudGripper (Zahid & Pokorny, 2023), Habitat 3.0 (Puig et al., 2023) and Causal World simulations (Ahmed et al., 2020; Bauer et al., 2022), greatly increasing the amount of available data. In addition, the maturity of physical robots and sensors also facilitates the generation of diverse multi-modal data. These real-world datasets represent different aspects of the environment. Moreover, sources like YouTube provide a wealth of information in the form of videos or general human behaviors.
2.2 Towards Foundation Veridical World Models
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption Action: Move forward, Look right, Open drawer Action: Pick cup, Put in drawer Action: Move forward, Look right, Pick cup Action: Put in drawer Task: Put the white cup in the drawer
Figure 2: A veridical world model can predict the consequences of sequence of actions and corresponding counterfactual correctly with confidence, ignoring irrelevant information like color of the cup. For example, the model should correctly predict that the cup will break if ‘open drawer’ action is not taken in the sequence. Images taken from AI2THOR (Kolve et al., 2017) simulator.
We believe that world models hold great promise for the future of embodied AI. One ideal world model is to facilitate the entire pipeline, spanning various tasks and platforms as shown in Figure 1. To achieve that, a desirable property is the ability to consume multi-modal inputs, generalize across environments, domains and tasks (i.e., being foundational). This requires the model to faithfully understand and model the world dynamics, i.e., being veridical. It needs to go beyond merely summarizing the content given what is observed, but also truly comprehend the world in a way that enables actionable interventions.
Combining the above two key aspects leads to the idea of FVWMs, or simply Foundation World Models (FWMs): a single (or set of) veridical, multi-modal models that can:
  1. (i)
    (Representation) Conceptually understand the components, structures, and interaction dynamics within a given system at different levels of abstractions (Olsen & Tylén, 2023);
  2. (ii)
    (Veridicality) Quantitatively model the underlying laws of such a system, that enables accurate predictions of counterfactual consequences of interventions/actions;
  3. (iii)
    (Foundational) Being able to generalize (i) and (ii) across diverse systems or domains encountered in the world.
When only (i) and (ii) are satisfied, FVWMs reduces to world models, originally intended to simulate possible futures of a single system (Ha & Schmidhuber, 2018; Li et al., 2020). By definition, FVWMs require a flexible formulation that can, using observational and experimental data from the world, reveal the underlying principles. This naturally falls into the realm of causality. However, incorporating such reasoning abilities while being able to process and generalize across multi-modal datasets and tasks is challenging, indicating substantial research opportunities for significant real-world impact.
3 Causality in Foundation World Models
In this section, we will demystify the next direction in causal research to be used for FVWMs. Specifically, we will: (i) briefly review SEMs, one of the commonly used canonical causal approaches; (ii) discuss the limitations of these canonical approaches in the context of FVWMs; (iii) describe the necessities of identifying new paradigms of causality research for FVWMs by demystifying a few common misconceptions.
3.1 Canonical approaches to causal modelling
Historically, causal research has been compartmentalized into distinct tasks. Causal discovery, for example, aims to recover causal relationships among variables from a dataset (Glymour et al., 2019; Spirtes et al., 2000), while causal inference is concerned with quantifying the effects of interventions. The advent of causal machine learning has expanded these horizons, introducing approaches such as end-to-end causal inference (Geffner et al., 2022), alongside advancements in causal representation learning (Schölkopf et al., 2021) and causal reinforcement learning (Zeng et al., 2023). There are two main frameworks to describe causality: potential outcomes (PO (Rubin, 2005; Hernán & Robins, 2010; Imbens & Rubin, 2015)) and Structural Equation Models (SEMs) (Pearl, 2009b; Peters et al., 2017). In the following, we briefly introduce the concept of SEM.
Structural Equation Models (SEM)
SEMs describe the generating mechanisms between variables. This typically follows the form of
$X_{i}=f_{X_{i}}(\text{Pa}_{X_{i}},U_{X_{i}}),$
where $f_{X_{i}}$ is a function mapping the values of exogenous variables $\mathcal{U}_{X_{i}}$ and its causes $\text{Pa}_{X_{i}}$ to the observation $X_{i}$.
A fully specified SEM generates a range of distributions that are systematically organized within the Pearl Causal Hierarchy (PCH) (Pearl et al., 2000), which categorizes causal information into three distinct levels: observational, interventional, and counterfactual. These levels reflect an ascending order of causal depth, each enabling a richer understanding of causality.
  1. (i)
    The observational level pertains to the distribution of observation variables without any interventions, typically represented as $p(\mathcal{X})$. It provides insights into natural associations and correlations.
  2. (ii)
    The interventional level involves altering the system through specific actions, indicated by Pearl’s $do$ operator. The interventional distribution, $p(\mathcal{X}_{-i}\,|\,do(X_{i}=a))$, characterizes the consequence of the action.
  3. (iii)
    The counterfactual level delves into hypothetical scenarios, considering what would occur had a different action been taken, given the same initial circumstances, expressed as $p(X_{j}\,|\,do(X_{i}=a),\mathcal{X}_{obs},\mathcal{U}_{obs})$.
The PCH provides a framework for reasoning in a causal-aware manner, allowing a model not only to observe the world but to interact with it as well.
Theory-oriented research paradigm
Given the difficulty to access the true system, theory-oriented research prevails in causality. For example, identifiability theory concerns the theoretical feasibility of deducing causal relations, interventional or counterfactual effects from observational data, assuming a certain set of models. Once identifiability is determined, the next steps involve developing methods for identifying (Pearl, 2009a; Angrist et al., 1996) and accurately estimating these causal quantities (Rosenbaum & Rubin, 1983; Stuart, 2010; Li et al., 2018; Chernozhukov et al., 2018; Wager & Athey, 2018). This theory-driven approach is one way to ensure the precision and trustworthiness of the model, as it contributes significantly to our comprehension of underlying causal mechanisms and guides decision making in relevant domains. However, this approach may also limit the method development considering real-world impact (Imbens, 2020).
Limitations
The canonical theory-oriented research has facilitated significant progress. Meanwhile, we are at a historical juncture to rethink this strategy for FVWMs and embodied AI due to the following limitations. Firstly, real-world sensory inputs are usually not structured into well-defined variables that admit a causal explanation (in terms of SEM/PO) and meaningful interventions (Schölkopf et al., 2021). Secondly, in many real-world scenarios, certain key theoretical properties, such as identifiability, may not hold, and their associated assumptions are often not verifiable (Pearl, 2009b; Robins, 2000). Thirdly, in the context of FVWMs, the case-specific nature of causal identification complicates the integration of causal reasoning (Bareinboim et al., 2022). Foundation models aim to provide a generalizable understanding of the world, but the bespoke nature of causal identification, which often relies on domain-specific assumptions, poses a significant barrier for the general environment. This discordance reveals a gap between the theoretical guarantee of causal reasoning and the practical needs of causal principles in foundation models, necessitating a re-evaluation of the approaches used to embed causality for FVWMs and embodied AI.
3.2 Demystifying causal modeling for FVWMs
The canonical causal research paradigm has inadvertently contributed to certain misconceptions regarding the field of causal machine learning, thus limiting its broader impact, especially under FVWMs and embodied AI. In the following, we aim to demystify three main misconceptions: (1) casual research is a separate field from Embodied AI and FVWMs; (2) causal research must be theory-oriented; and (3) predictive and causal models are different.
Misconception I: Causal ML is a separate field from Embodied AI and FVWMs.
Despite causality being pivotal for the generalization of world models (Anonymous, 2024b), Causal ML is often overlooked in discussions, compared to RL and foundation models. This oversight partly stems from distinct tools, such as SEMs and PO, and application domains traditionally associated with canonical causal research. However, it is crucial to recognize that Causal ML is intricately connected to both reinforcement learning (bandit and imitation learning as a special case) and foundation models with shared goals.
Bandits and RL share the same goal of learning the best action policy from observational data. Offline RL fundamentally shares the same goal as causal modeling, while online RL focuses more on learning from interactions (Kaddour et al., 2022). While RL does not always explicitly model environmental dynamics – as seen in model-free RL – it inherently captures these dynamics within the learned policy or value function through interactions with environment. Thus, RL is intrinsically causal in a board sense considering that it need to inherently understand the consequence of actions for policy learning. A more explicit connection can be seen in model-based RL, where temporal causal models can be used to formulate the transition dynamics. Such temporal causal models induce the transition probabilities between states across time. Actions by agents are then viewed as interventions to this model. Leveraging ideas from causality would allow the learned dynamics to only reason over causally relevant attributes such as force and location to predict the next state (e.g. location of the cup in Figure 2). Many recent works have begun to incorporate causal concepts into the state transition dynamics (see Section 6). This potentially enhances sample efficiency and generalizability beyond training contexts.
Large multi-modal foundation models aim to facilitate reasoning and planning (Sun et al., 2023; Huang & Chang, 2022; Webb et al., 2023; Li et al., 2023), which inherently requires understanding the physical nature of environments. Although much progress has been made (Wei et al., 2022; Zhang et al., 2022; Creswell et al., 2022; Wang et al., 2023a; Green et al., 2012), they are trained to gather information only at the observational level, which may limit their understanding of the physical world (lack of interventional and counterfactual information). While these models, including CLIP (Radford et al., 2021) and GPT-4 (Achiam et al., 2023), excel in various tasks like vision-language understanding and image recognition with multi-modal input, they are limited by issues like hallucination (Huang et al., 2023a; Gunjal et al., 2023) and reduced multi-step reasoning abilities (Jin et al., 2024; Sun et al., 2023), indicating the lack of causal understanding. On the other hand, the canonical causality research fails to incorporate multi-modal dataset and is typically tied with predefined variables for SEM. Recent work (discussed in Section 6) reveals the synergy between Causal ML and foundation models for mutual enhancement, and paves the way for FVWMs.
Misconception II: Causal ML research must be theory-oriented.
The traditional emphasis of theory-oriented approach in causal machine learning research often overshadows the empirical-driven approach, that prioritizes the causal experimentation, verification, and hypothesis selection (Dorie et al., 2019; Jensen, 2019; Gong et al., 2022; Nie & Wager, 2021; Saito & Yasui, 2020; Vansteelandt et al., 2012; Schuler et al., 2018; Rolling & Yang, 2014; Matthieu & Gaël, 2023; Cook et al., 2002; Hoover, 2001; Bareinboim & Pearl, 2016; Bareinboim, 2014). Consequently, over-reliance on the pure theoretical approach can lead to models that fail to perform effectively in practical, complex scenarios. We would like to emphasize that one of the end goals of causality research is causal-aware reasoning and to predict the consequences accurately given any actions. Canonical formulations with theoretical guarantee is one way for causal research, while focusing on evaluating the causal property and empirical results is an alternative when it comes to FVWMs. Such an empirical-driven shift allows models to assimilate causal knowledge from a wide array of scenarios, akin to the way large language models are evaluated. Moreover, the unique problem setting of embodied AI opens up a realm where empirical evaluation can become the cornerstone of causal model assessment. Through interaction with both real and simulated environments, embodied AI agents can generate and leverage interventional or counterfactual data, thereby permitting a more pragmatic approach to causal understanding. Thus, the empirical-driven causal research under FVWMs and embodied AI is equally or more important than the theory-oriented paradigm.
Misconception III: Predictive models cannot be causal.
The commonly understood principle that correlation does not imply causation is particularly relevant in the field of predictive modeling, and it is often misunderstood as predictive models cannot be causal. This interpretation holds some truth when models are trained solely on observational data, which are prone to confounding effects that can bias predictions about the outcomes of specific actions. However, this changes with access to active intervention data, allowing the predictive model to demonstrate causal properties. This aligns with the discussion in misconception I : RL models are inherently causal and can learn to take optimal actions for a variety of tasks (Kaddour et al., 2022; Anonymous, 2024b). Thus, causal models are inherently high-accuracy prediction models of consequences about all possible actions. Predictive models can also be considered causal models when trained with appropriate algorithms and datasets, even without explicit causal theory considerations. In the realm of foundation models for embodied AI, greater opportunities arise from the potential interaction with the physical world. This allows the integration of both causal considerations and large-scale prediction models (Ke et al., 2019; Scherrer et al., 2021; Toth et al., 2022).
4 Research opportunities
Recognizing the immense potential of foundation models for embodied AI and the necessity of a causal perspective to ensure accurate estimation of action consequences, we foresee significant research opportunities, especially since current foundation models are not causal-aware. In particular, we identify a pressing need for research advancement in the following aspects.
Diverse modalities
The foundation model for Embodied AI extends beyond traditional vision and language realms. Robotics necessitates the integration and interpretation of diverse sensory data, including tactile and various torque sensors, covering a broad range of modalities which vary from one agent to the next. This requirement surpasses the capabilities of current vision and language-based foundation models. Developing a system capable of effectively processing and utilizing these complex sensory data is crucial for enhancing robotic capabilities, especially in more intricate and nuanced tasks. On the other hand, building a causally-aware FWMs that comprehends high-dimensional observations from a large number of potential different modalities to model the underlying mechanisms is very challenging and presents multiple research opportunities:
First, integrating data from multiple modalities to achieve a cohesive understanding from different sensory inputs poses challenges in spatial and temporal alignment (Ward-Cherrier et al., 2018; Lin et al., 2023b; Wang et al., 2023b; Huang et al., 2023b; Behley et al., 2019) due to the varying perspectives and information overlap provided by each sensor, as well as the need to reconcile different sampling intervals to maintain an up-to-date representation of the world.
Secondly, agents vary widely in their sensor and effector configurations, from humanoid robots like Tesla’s Optimus to quadrupedal robots like Boston Dynamics’ Spot, to virtual cars, each with unique sensory inputs and available effectors. For FWMs to make precise action-outcome predictions, they must semantically interpret these diverse inputs and controls, necessitating a representational model that encapsulates the essence of an agent’s design to enable generalization across different configurations.
Lastly, handling the complexities of varying abstraction levels for effective interaction presents a formidable challenge (Schölkopf et al., 2021; Wang & Jordan, 2021; Koul et al., 2023; Zhang et al., 2020c; Lamb et al., 2022). In standard causal models, variables are precisely defined within structural equation models (SEMs) for a single dataset and task. However, such clarity is often absent in the realm of multi-modal, high-dimensional inputs across diverse tasks. While emerging works are exploring causal-aware foundation models (Zhang et al., 2023c; Ke et al., 2022), they remain confined to tabular data with predefined tasks. Given a high-dimensional input, a representation meaningful for interactions and their effects (affordances) is necessary to ensure the world model is meaningful. For instance, modeling interaction and consequences at a pixel level provides little immediate insight and requires additional transformations to either map the (motor) control of an agent into actions in pixel space or to extract an interpretable understanding of the effect the interaction had. Additionally, a foundational world model for embodied AI, which aims to achieve broad generalization, must consider different knowledge hierarchies to ensure effective generalization and knowledge transfer across varied contexts. For example, understanding the consequences of avoiding a car attempting to overtake should be consistent regardless of the specific car instances (car, brand), and the same principle applies to home robotics with different effectors for similar tasks. The FVWM, therefore, requires hierarchical representations that facilitate knowledge transfer at multiple levels, enabling nuanced understanding and application across various scenarios and tasks.
New paradigm for interactions
As discussed, the canonical research in causality and RL shares the same objective but primarily differs in the main source of the data paradigm: observational data and intervention data. Traditionally, causality seeks to learn the world model using existing data, assuming that new interventions are not feasible, while RL learns through interactions. For embodied AI with FVWM, there will be access to a vast amount of offline data from humans, such as vision language data, human motion data, and robotics data, along with the possibility to obtain new intervention data. We need to utilize all different data paradigms in the most efficient way, breaking down the barriers between foundation models, RL, and causality. For example, we can look into debiasing the observational data using causal-aware methods such as CInA (Zhang et al., 2023c). Moreover, how to efficiently combine these algorithms and data paradigms in a single foundation model is also non-trivial with a lot of potentials. For example, if the model has a high loss of predicting how the dynamics change when a modality in the past is altered, then it should aim to answer this either through weighting the loss higher for the altered data point, or aim to collect this information through online interactions. This can be done in an iterative manner where a base model is trained from offline data, then allowed access to the online simulator to satisfy its curiosity (i.e. counterfactuals) and so on and so forth (Zeng et al., 2023; Gasse et al., 2021; Ivanova et al., 2023; Tigas et al., 2022; Seitzer et al., 2021; Sontakke et al., 2021). Such causally curious foundation model can improve the sample efficiency of training large models by reducing computing times, enabling multi-step reasoning and reduced hallucinations, and thereby no longer requiring RLHF (Casper et al., 2023) to finetune the trained models for specific tasks. A veridical causal foundation model can therefore be directly employed to make multi-step future predictions and directly enable planning methods (Zhang et al., 2023d) to find the sequence of the path leading to the desired outcomes. This is also consistent with the consideration of active data collection, where RL and experimental design are closely interlinked (González et al., 2016; Ma et al., 2019; Yin et al., 2020; Foster et al., 2021). Additionally, such an FVWM can massively accelerate planning methods to learn optimal policies in a sample efficient manner.
Empirical driven research and evaluation
In incorporating causality research into the foundation models for embodied AI, the key lies in the successful execution of actions, which depends on accurately predicting or imagining the consequences of different actions. In embodied AI research, physical meaningfulness is essential. Therefore, we believe that shifting the research culture towards being empirically driven, especially for causal researchers, and developing relatable, realistic, and economical evaluation methods is crucial for success.
Evaluating real-world scenarios with a physical device is very costly in terms of time and potential risks, although it is necessary for the ultimate success of foundation embodied AI. Thus, constructing a curriculum of tasks that are easy to evaluate is key. This means evaluations can be performed at scale, replicated by any research institute, have clear metrics demonstrating progress, and provide a reliable approximation for end tasks in the real world. This requires not only further advancements in embodied AI in simulated environments like CausalWorld (Ahmed et al., 2020; Bauer et al., 2022) and Habitat 3.0 (Puig et al., 2023) but also the creation of innovative tasks that aid in this process. For example, evaluation on counterfactual prediction can be a good approximation for assessing true world understanding. However, we need to consider concrete tasks and evaluation metrics. In this case, the task could involve a generated video (Yin et al., 2023), while the metric needs to be more nuanced, as pixel space evaluation may have drawbacks, as discussed in computer vision research communities.
Moreover, it’s unlikely that a single metric will fully capture the model’s performance, especially with the new causal perspective. Traditional metrics such as graph prediction accuracy or treatment effect estimation accuracy with well-defined variables become less relevant. However, the philosophical approach and perspective in evaluating causal properties remain crucial. For example, it’s worth considering if evaluating causal metrics in a representation space will provide additional insights complementary to prediction tasks. Also, the key to embodied AI is understanding affordances and the consequence of actions, so evaluating counterfactuals in an imagined world might be effective.
On the other hand, the primary reason that correlation-based models are often considered unsuitable for causal-aware tasks is the bias present in the datasets collected. A causal-aware model is expected to accurately predict future outcomes in out-of-distribution evaluations (similar to A/B testing data), even when trained with biased data. Therefore, it’s crucial to consider the availability of training data and the types of properties that reflect real-world scenarios, while also ensuring that the evaluation dataset genuinely represents generalization across various scenarios.
In conclusion, while the broader concept of embodied AI encompasses virtual and mixed reality environments, as well as physical real-world environments (robotics and human action) , the final evaluation, especially considering human safety, cannot be overlooked (Kadian et al., 2020; Ma & Zhang, 2023; Jin* et al., 2023; Yang et al., 2023). All these efficient evaluation methods aim to pave the way for real-world deployment.
5 Deploying Embodied AI
Many factors affect how FVWMs can impact the deployment of Embodied AI in the real world. In this chapter, we focus on what we believe will be one of the most important factors, which is FVWMs ability to greatly scale up the deployments of Embodied AI in robotics, and thereby also robotics in general.
General purpose robots
General purpose robots form perhaps the most exciting prospect of generally capable Embodied AI. By compatibility with human environments and tools, robots of this kind could finally bridge the gap between the physical and digital worlds and allow large scale automation of physical tasks. Potential applications range from automating domestic tasks (Puig et al., 2023), nursing (Ohneberg et al., 2023), work in hazardous environments (Petereit et al., 2019) to manufacturing (Kheddar et al., 2019). We discuss ethical considerations regarding societal impact in our impact statement. Recent work has demonstrated zero-shot solutions to complex tasks, yet shows poor performance when faced with new environments. Research in causal FVWMs should prioritize consideration of diverse modalities and platforms, enabling success across a wide range of tasks and environments.
Rapid deployment of specialized robots
While general purpose robots capable of zero-shot learning stand to reduce deployment time, specialized robots will remain more cost effective in some scenarios - especially for large-volume production. Examples include farming (Aravind et al., 2017), scientific lab exploration (Ren et al., 2023; Boiko et al., 2023) and industrial welding machines (Grau et al., 2020). Implementing specialized robots is a time-consuming process and with the learning to perform task without explicit programming is prioritized on the research agenda (Dzedzickis et al., 2022). However, current learning implementations are typically limited to a single environment (Meyes et al., 2017). We believe that causal-aware FVWMs may solve these constraints and can therefore dramatically impact the deployment of industrial specialized robots. An important testbed for this work will therefore be operation across robot morphologies specialized with different sensing and tooling attachments.
Robustness and safety
Industrial robots have typically been confined to isolated cells. Through constraints such as force limitations, more collaborative modes of operation have been enabled, which have resulted in improvements in productivity (Dzedzickis et al., 2022). In other applications, such as nursing homes, direct physical interaction with humans is vital. Yet in such situations, the constraints from highly controlled industrial environments may not be sufficient to ensure physical safety. A thorough grasp of the outcomes resulting from physical interactions, facilitated by FVWMs, presents an opportunity to enhance deployment safety, provided that the evaluation steps outlined in Section 4 are meticulously followed.
6 Related Work
Foundation model for Embodied AI
There have been a number of surveys on foundation models with a focus on recent research advances, challenges, and opportunities (Zhou et al., 2023a; Bommasani et al., 2021). Regarding embodied AI, Firoozi et al. (2023); Xiao et al. (2023) survey the role of pretrained foundation models in robotics for perception, navigation, decision making and control. Recent studies have explored the use of vision-language FMs in embodied AI tasks to improve sample efficiency and performance abilities (Khandelwal et al., 2022; Batra et al., 2020; Shridhar et al., 2022). Moreover, numerous surveys from the perspective of robotic tasks have been published, providing insights applicable to both deep learning and machine learning at large (Yin et al., 2021; Newbury et al., 2023). Recent research advances have increasingly focused on utilizing foundation models for embodied AI. A number of methods leverage large LLMs and VLMs for planning, re-planning and navigation (Lin et al., 2023a; Schumann et al., 2023; Skreta et al., 2024). Also, more recent work focuses on generalization across different environments and tasks for robots based on VLM, such as SayCan (Brohan et al., 2023b), RT-2 (Brohan et al., 2023a), GATO (Reed et al., 2022), AutoRT (Ahn et al., 2024), and autonomous driving, such as GAIA (Hu et al., 2023).
Causality and Causal Foundation Model
In recent years, books (Pearl, 2009b; Hernán & Robins, 2010; Cunningham, 2021; Imbens & Rubin, 2015; Spirtes et al., 2000; Peters et al., 2017) as well as surveys (Spirtes & Zhang, 2016; Kitson et al., 2021; Shojaie & Fox, 2022; Zanga et al., 2022) have been published, drawing attention to causality and causal machine learning in general. Related to our work is the recent overview of open problems from (Kaddour et al., 2022), which connect different machine learning approaches to causality (Schölkopf, 2022); and a survey and position paper on causal reinforcement learning (Zeng et al., 2023) which aligns with our position of breaking the boundary of these field for advanced performance. There is large body of research work considering generic deep learning and causality including (Schölkopf et al., 2021; Jin et al., 2022; Zhou et al., 2023b; Berrevoets et al., 2023).
With the advent of foundation models, particularly large language models (LLMs), there have been multiple position and evaluation papers considering whether LLMs are causally aware and if they can be used to complete causal tasks (Willig et al., 2022; Tu et al., 2023; Kıcıman et al., 2023; Zhang et al., 2023a, b; Jin et al., 2023; Jin* et al., 2023; Zečević et al., 2023). Generally, these papers argue that existing LLMs can discuss causality but are not capable of reasoning in a causally aware manner. Consequently, this further confirms that current LLMs are not yet world models. Meanwhile, recent research has begun to show that it may be possible to innovate based on transformer-style architecture (Lorch et al., 2022; Zhang et al., 2023c), enabling foundation models to estimate treatment effects. Although these works are in their early stages, they provide guidance and the possibility for a world model that is causally aware, which will be essential for embodied AI.
Causal World Models
World models, often used in RL, aim to model the transition dynamics which are expected to be intrinsically causal. It is a longstanding challenge, where early work has used standard sequence models for this purpose. For example, Ha & Schmidhuber (2018) employ a Variational Autoencoder (VAE) to condense visual inputs into a latent representation with a RNN to model the dynamics. Other studies have focused on learning latent state dynamics models, as seen in Hafner et al. (2019); Gelada et al. (2019), while advancements in the Dreamer series (Hafner et al., 2020, 2023), TransDreamer (Chen et al., 2022) and Micheli et al. (2022) highlight the benefits of discrete latent variables and transformers for latent dynamics. Recent works have also leveraged diffusion models directly for world models (Rigter et al., 2023; Anonymous, 2024a).
The community has increasingly recognized the important role of causal properties in world models. Recently, Anonymous (2024b) proved in theory that any agents capable of generalizing under domain shift must learn a causal model, and there have been many works trying to incorporate causality. Zhang et al. (2020b, 2019) use causal core sets (Zhang et al., 2020a) to model latent transition dynamics. Li et al. (2020); Lu et al. (2018), on the other hand, considers the confounding effect during the modelling of state transitions. Another line of work (Annabi, 2022; Volodin et al., 2020; Zhu et al., 2022; Wang et al., 2022; Tomar et al., 2021) explicitly build a SEM among variables in the states, allowing the transition dynamics to only depend on causally-relevant feature. Unfortunately, existing causal world models fall short as FWMs due to their inability to generalize across domains and tasks, and to process multi-modal inputs efficiently. Future research on causal-aware FVWM will bridge the gap and enable embodied AI.
7 Conclusion
In this position paper, we advocate for FVWMs as key to unlocking the full potential of embodied AI in both general and scalable contexts, highlighting the indispensability of causality in these models. We reiterate that causality as a property is fundamental to accurate action and consequence prediction, and its integration is crucial in the realm of machine learning and foundation models. Looking forward, we envision future research focusing not only on high-modality observations but also on transcending the boundaries of various machine learning sub-fields. This approach aims to optimize data utilization through both physical interactions and a blend of online and offline methodologies. Moreover, we emphasize the significance of establishing empirical evaluation benchmarks to foster real-world impact in a cost-effective and low-risk manner.
While this paper concentrates on the FVWMs in embodied AI, we acknowledge that numerous other research challenges remain unaddressed by this study. These include, but are not limited to, advancements in hardware, human-robot communication and collaboration, and comprehensive socio-economic considerations, which are very important for embodied AI in general.
Impact statement
This position paper, focusing on world models for Embodied AI, does not propose new technological advancements with direct societal consequences in isolation. However, it is crucial to address the broader social implications of the field. This is a topic of intense debate with a multitude of works predicting its outcome and potential mitigation. However, a reoccurring theme in the debate is that the whole of society to ensure an equitable transition to rapidly increasing automation. Willcocks (2020) argues that such a transition will unlikely result in a net job loss, and will be vital to mitigate changing demographics which are predicted to lead to labour shortages. Using foundation models in embodied AI shifts the focus of automation from tasks that are dirty, dull, and dangerous (3D) to those that necessitate advanced cognitive and reasoning skills. As a result, the implications of widespread deployment in automating various job types could vary significantly, necessitating careful consideration by society.
Another important consideration is that Embodied AI should be ethical and fair. For the real-world evaluation of embodied AI, safety and ethical considerations are key, as there is the possibility of inducing physical harm. Moreover, Causal FVWMs may reduce biases and lead to fairer algorithms (Kusner et al., 2017; Zhang & Bareinboim, 2018). However, as suggested by Nilforoshan et al. (2022) may not directly translate to equitable decision making. Instead it will remain important for researchers to develop this area further.
Acknowledgment
The authors wish to express gratitude to Katja Hofmann, Nan Duan, Anastasia Varava, Xin Tong, Tom Minka, Jiang Bian, David Sweeney and Baining Guo, especially to John Langford, whose discussions have inspired and aided the manuscript.