### v2

# The Essential Role of Causality in Foundation World Models for Embodied AI

Abstract
Recent advances in foundation models, especially in large multi-modal models and conversational agents, have ignited interest in the potential of generally capable embodied agents. Such agents will require the ability to perform new tasks in many different real-world environments. However, current foundation models fail to accurately model physical interactions and are therefore insufficient for Embodied AI. The study of causality lends itself to the construction of veridical world models, which are crucial for accurately predicting the outcomes of possible interactions. This paper focuses on the prospects of building foundation world models for the upcoming generation of embodied agents and presents a novel viewpoint on the significance of causality within these. We posit that integrating causal considerations is vital to facilitating meaningful physical interactions with the world. Finally, we demystify misconceptions about causality in this context and present our outlook for future research.
Machine Learning, ICML
1 Introduction
The ability to perform physically meaningful interactions in real-world environments is essential to maximizing the potential of technology in everyday life. Entities capable of conducting interactions of this kind are in our work referred to as embodied agents. These agents can vary in form, ranging from mechanical robots to humans augmented by AI in mixed or virtual reality environments where interactions bear physical importance.
Foundation models account for a significant advancement in generalization across a range of tasks with multi-modal input, compared to canonical deep learning models from the past decade. A single foundation model can address a wide variety of tasks in a zero-shot manner. For example, in robotics, we anticipate that foundation models for embodied agents will make the introduction of new robots into our life effortless: each new robot can immediately solve a given task using the same model without the need for a specialized model. In addition, robots can quickly adapt to a new environment by interacting with it and using the new experiences to improve without lengthy retraining or complex programming.
However, contemporary foundation models are not sufficient. Current approaches, dominated by large (vision-) language models (Achiam et al., 2023; Bubeck et al., 2023; Chen et al., 2023), are based on correlational statistics and do not explicitly capture the underlying dynamics, compositional structure or causal hierachies. The lack of a veridical world model renders them unsuitable for use in Embodied AI, which demands precise or longterm action planning, efficient and safe exploration of new environments or quick adaptation to feedback and the actions of other agents.
World models (Ha & Schmidhuber, 2018) make action-conditioned predictions about the future state of the environment by training on action-labelled interaction data. World models therefore capture the consequences of actions and are suitable for optimising decision-making (Sutton, 1991; Janner et al., 2019). To train general world models it will be necessary to leverage large readily-available observational datasets, such as internet videos, in addition to action-labelled data which is much more scarce and expensive to generate. The best way to utilise observational data for world modeling remains an open question (Seo et al., 2022; Bruce et al., 2024). In this paper we argue that causal considerations are necessary to unlock the potential for training world models from internet-scale data.
Causality at its core aims to understand the consequences of actions, allowing for interaction planning. Philosophy and cognitive sciences sees understanding causal concepts as both the basics and sometimes the ultimate goal for humans to learn how to interact with the world (Gibson, 1978; Gopnik et al., 2007; Adams & Aizawa, 2021) and key in children’s development (Piaget, 1965; Gibson, 1988). This aligns with the concept of affordances in Embodied AI research (Kjellström et al., 2011; Koppula et al., 2013; Koppula & Saxena, 2013; Ardón et al., 2020; Ahn et al., 2022), which pertains to the types of action that can be applied and the resulting consequences. Importantly, even with the help of available real or simulated environments, the experimentation might still be too coarse-grained to deal with spurious relationships (Herd & Miles, 2019; Lavin et al., 2021). Thus, the world model for an agent, including humans, should be causal-aware (Richens & Everitt, 2024) and able to efficiently incorporate data from observations, demonstrations, and interaction (i.e., the property referred as veridicality in this paper). As such, causality offers tools and insights that hold the key pieces to building Foundation Veridical World Models (FVWMs) (Section 2.2) that will power future embodied agents.
In this position paper, we argue that future progress in Embodied AI relies on further advances in foundation models (Section 2). For such research, causality is pivotal in developing the FVWM, which requires us to go beyond canonical causal research paradigms (Section 3). Thus, we foresee research opportunities (Section 4) advancing foundation world models in Embodied AI from a causal perspective. We extend our discussion with considerations for the eventual deployment of Embodied AI (Section 5).
2 Embodied AI and Foundation Veridical World Models
Foundation Veridical World Model Daily assistance humanoid Industrial manufacturing robots Mixed-reality interactions … …
Figure 1: The Foundation Veridical World Model (FVWM) is designed to comprehend associations, counterfactuals, and interactions within the world, aiding any embodied AI agent in executing tasks across diverse environments and platforms.
For many years, research has pursued the creation of Embodied AI. However, existing approaches are limited to performing narrow tasks in highly constrained environments, such as assembling machines in production lines. This inhibits the applicability of existing Embodied AI systems to general real-world settings, where environments are often dynamic and unstructured, and the set of tasks to be performed is open-ended. This highlights a significant gap between current technologies and fulfilling the promises of general Embodied AI. We believe that recent technological advancements (Section 2.1), including hardware, foundation models, and data, have brought general Embodied AI within reach, where the remaining gap can be overcome by veridical world models (Section 2.2).
2.1 The readiness of Embodied AI research
Hardware advancement
The pace of innovation in hardware and systems, ranging from real-world robots to virtual and mixed reality devices, is advancing at an unprecedented rate, though there is still a significant journey ahead before achieving maturity for general tasks. The costs of both humanoid (e.g. Optimus, Phoenix (Sanctuary, 2024), Digit (Agility, 2024)) and quadrupedal robots (e.g. Spot (BostonDynamics, 2024), Go2 (Unitree, 2024) ) have been rapidly decreasing over the last decade, with a huge potential market size (GoldmanSachs, 2023; Tilley, 2017). For virtual mixed reality experiences, Oculus from Meta has increased in popularity and Apple’s Vision Pro is taking a commercial leap into offering new interactive experiences, creating opportunities for virtual agents. In addition, new types of sensors for robots have been invented recently. Electronic skin can capture the sensation of touch (Sanderson, 2021) to enable dexterous robots (Lambeta et al., 2020; Sun et al., 2022), while perovskite retinomorphic sensors (Trujillo Herrera & Labram, 2020) mimic the human retina and react to changes in illumination rather than constant signals. These advancements offer different ways to sense the world. This enables the consideration of more dextrous manipulation tasks, such as those involving deformable objects, rather than rigid ones (Yin et al., 2021).
Meanwhile, the evolution of AI computing power has accelerated at an unprecedented rate. This surge in computing capabilities is primarily driven by innovations in hardware, such as more powerful GPUs and specialized AI processors (Dally et al., 2021; Saravanan & Kouzani, 2023), providing the computing basis for foundation models. These developments have enabled AI systems to process vast amounts of data more quickly and more effectively than ever before, expanding the horizons of potential across various domains.
Paradigm shift in AI
The paradigm shift driven by foundation models in both language and vision has markedly enhanced the ability of AI models to understand and perceive the environment (Bubeck et al., 2023; Achiam et al., 2023; Yuan et al., 2021; Liu et al., 2023; Chen et al., 2023). These systems now demonstrate an increasing capability to generalize on high-level planning and decision-making tasks, often in a zero-shot manner (Ahn et al., 2022). This has led to a shift in the research paradigm from specialized models for a single target task and dataset to foundation models generalizing across many tasks and datasets. These increased generalization capabilities hold the key to building world models suitable for general Embodied AI systems.
Data availability
Adding to the timeliness of this venture is the increasing amount of of data in the field of Embodied AI, which is crucial for effective training of foundation world models. Robotics simulation environments and demonstration datasets have become more diverse, encompassing different types of tasks and scenarios, such as OpenX embodiment (Open X-Embodiment Collaboration et al., 2023), Isaac gym (Makoviychuk et al., 2021; Ma et al., 2023), CloudGripper (Zahid & Pokorny, 2023), Habitat 3.0 (Puig et al., 2023) and Causal World simulations (Ahmed et al., 2020; Bauer et al., 2022), greatly increasing the amount of available data. In addition, the maturity of physical robots and sensors also facilitates the generation of diverse multi-modal data. These real-world datasets represent different aspects of the environment. Moreover, sources like YouTube provide a wealth of information in the form of videos or general human behaviors (Seo et al., 2022).
2.2 Towards Foundation Veridical World Models
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption Action: Move forward, Look right, Open drawer Action: Pick cup, Put in drawer Action: Move forward, Look right, Pick cup Action: Put in drawer Task: Put the white cup in the drawer
Figure 2: A veridical world model can predict the consequences of a sequence of actions and corresponding counterfactuals correctly with confidence, ignoring irrelevant information like color of the cup. For example, the model should correctly predict that the cup will break if ‘open drawer’ action is not taken in the sequence. Images taken from AI2THOR (Kolve et al., 2017) simulator.
We believe that world models hold great promise for the future of Embodied AI. Ideally, a general world model will facilitate the entire Embodied AI pipeline, spanning various tasks and platforms as shown in Figure 1. To achieve that, a required property is the ability to consume multi-modal inputs, and generalize across environments, domains, and tasks (i.e., being foundational). This requires the model to faithfully understand and model the world dynamics, i.e., being veridical. It needs to go beyond merely summarizing the content given what is observed, but also truly comprehend the world in a way that enables actionable interventions.
Combining the above two key aspects leads to the idea of FVWMs, or simply Foundation World Models (FWMs), a single (or set of) veridical, multi-modal models that can:
  1. (i)
    (Representation) Conceptually understand the components, structures, and interaction dynamics within a given system at different levels of abstractions (Olsen & Tylén, 2023);
  2. (ii)
    (Veridicality) Quantitatively model the underlying laws of such a system, that enables accurate predictions of counterfactual consequences of interventions/actions;
  3. (iii)
    (Foundational) Generalize (i) and (ii) across diverse systems or domains encountered in the world.
When only (i) and (ii) are satisfied, FVWMs reduces to world models, originally intended to simulate possible futures of a single system (Ha & Schmidhuber, 2018; Li et al., 2020). To accomplish (iii), FVWMs require a flexible formulation that can, using observational and interaction data from diverse real-world datasets, reveal the underlying principles that generalize across environments. This naturally falls into the realm of causality. However, incorporating such causal reasoning abilities while being able to process and generalize across multi-modal datasets and tasks is challenging, indicating substantial research opportunities for significant real-world impact.
3 Causality in Foundation World Models
In this section, we will demystify the next direction in causal research to be used for FVWMs. Specifically, we will: (i) briefly review SEMs, one of the commonly used canonical causal approaches; (ii) discuss the limitations of these canonical approaches in the context of FVWMs; (iii) describe the necessities of identifying new paradigms of causality research for FVWMs by demystifying a few common misconceptions.
3.1 Canonical approaches to causal modelling
Historically, causal research has been compartmentalized into distinct tasks. Causal discovery, for example, aims to recover causal relationships among variables from a dataset (Glymour et al., 2019; Spirtes et al., 2000), while causal inference is concerned with quantifying the effects of interventions. The advent of causal machine learning has expanded these horizons, introducing approaches such as end-to-end causal inference (Geffner et al., 2022), alongside advancements in causal representation learning (Schölkopf et al., 2021) and causal reinforcement learning (Zeng et al., 2023). There are two main frameworks to describe causality: potential outcomes (PO (Rubin, 2005; Hernán & Robins, 2010; Imbens & Rubin, 2015)) and Structural Equation Models (SEMs) (Pearl, 2009b; Peters et al., 2017). In the following, we briefly introduce the concept of SEM.
Structural Equation Models (SEM)
SEMs describe the generating mechanisms between variables. This typically follows the form of
$X_{i}=f_{X_{i}}(\text{Pa}_{X_{i}},U_{X_{i}}),$
where $f_{X_{i}}$ is a function mapping the values of exogenous variables $\mathcal{U}_{X_{i}}$ and its causes $\text{Pa}_{X_{i}}$ to the observation $X_{i}$.
A fully specified SEM generates a range of distributions that are systematically organized within the Pearl Causal Hierarchy (PCH) (Pearl et al., 2000), which categorizes causal information into three distinct levels: observational, interventional, and counterfactual. These levels reflect an ascending order of causal depth, each enabling a richer understanding of causality.
  1. (i)
    The observational level pertains to the distribution of observation variables without any interventions, typically represented as $p(\mathcal{X})$. It provides insights into natural associations and correlations.
  2. (ii)
    The interventional level involves altering the system through specific actions, indicated by Pearl’s $do$ operator. The interventional distribution, $p(\mathcal{X}_{-i}\,|\,do(X_{i}=a))$, characterizes the consequence of the action.
  3. (iii)
    The counterfactual level delves into hypothetical scenarios, considering what would occur had a different action been taken, given the same initial circumstances, expressed as $p(X_{j}\,|\,do(X_{i}=a),\mathcal{X}_{obs},\mathcal{U}_{obs})$.
The PCH provides a framework for reasoning in a causal-aware manner, allowing a model not only to observe the world but to interact with it as well.
Theory-oriented research paradigm
Given the difficulty of accessing the true system, theory-oriented research prevails in causality. For example, identifiability theory concerns the theoretical feasibility of deducing causal relations, interventional or counterfactual effects from observational data, assuming a certain set of models. Once identifiability is determined, the next steps involve developing methods for identifying (Pearl, 2009a; Angrist et al., 1996) and accurately estimating these causal quantities (Rosenbaum & Rubin, 1983; Stuart, 2010; Li et al., 2018; Chernozhukov et al., 2018; Wager & Athey, 2018). This theory-driven approach is one way to ensure the precision and trustworthiness of the model, as it contributes significantly to our comprehension of underlying causal mechanisms and guides decision making in relevant domains. However, as we shall discuss, this approach may limit the real-world impact of methods developed (Imbens, 2020).
Limitations
The canonical theory-oriented research has facilitated significant progress. Meanwhile, we are at a historical juncture to rethink this strategy for FVWMs and embodied AI due to the following limitations. First, real-world sensory inputs are usually not structured into well-defined variables that admit a causal explanation (in terms of SEM/PO) and meaningful interventions (Schölkopf et al., 2021). Second, in many real-world scenarios, certain key theoretical properties, such as identifiability, may not hold, and their associated assumptions are often not verifiable (Pearl, 2009b; Robins, 2000). Third, in the context of FVWMs, the case-specific nature of causal identification complicates the integration of causal reasoning (Bareinboim et al., 2022). Foundation models aim to provide a generalizable understanding of the world, but the bespoke nature of causal identification, which often relies on domain-specific assumptions, poses a significant barrier for general environments. This discordance reveals a gap between the theoretical guarantee of causal reasoning and the practical needs of causal principles in foundation models, necessitating a re-evaluation of the approaches used to embed causality for FVWMs and embodied AI.
3.2 Demystifying causal modeling for FVWMs
The canonical causal research paradigm has inadvertently contributed to certain misconceptions regarding the field of causal machine learning, thus limiting its broader impact, especially in the context of FVWMs and embodied AI. In the following, we aim to demystify three main misconceptions: (1) casual research is a separate field from Embodied AI and FVWMs; (2) causal research must be theory-oriented; and (3) predictive and causal models are different.
Misconception I: Causal ML is a separate field from Embodied AI and FVWMs.
Despite causality being pivotal for the generalization of world models (Richens & Everitt, 2024), Causal ML is often overlooked in Embodied AI discussions, compared to RL and foundation models. This oversight partly stems from distinct tools, such as SEMs and PO, and the application domains traditionally associated with canonical causal research. However, it is crucial to recognize that Causal ML is intricately connected to both reinforcement learning (bandits and imitation learning as a special case) and foundation models, with shared goals.
Bandits and online RL share the aim of learning the best action policy from online interactions. Offline RL assumes access to only fixed dataset from the system of interest, and therefore considers the same problem setting as causal modeling (Kaddour et al., 2022). While RL does not always explicitly model environmental dynamics – as seen in model-free RL – it implicitly captures these dynamics within the learned policy or value function. Thus, RL is intrinsically causal in the sense that it needs to inherently understand the consequences of actions for policy learning. A more explicit connection can be seen in model-based RL, where temporal causal models can be used to formulate the transition dynamics. Such temporal causal models induce the transition probabilities between states across time. Actions by agents are then viewed as interventions to this model. Leveraging ideas from causality would allow the learned dynamics to only reason over causally relevant attributes such as force and location to predict the next state (e.g. location of the cup in Figure 2). Many recent works have begun to incorporate causal concepts into the state transition dynamics (see Section 6). This potentially enhances sample efficiency and generalizability beyond training contexts.
Large multi-modal foundation models aim to facilitate reasoning and planning (Sun et al., 2023; Huang & Chang, 2022; Webb et al., 2023; Li et al., 2023), which inherently requires understanding the physical nature of environments. Although much progress has been made (Wei et al., 2022; Zhang et al., 2022; Creswell et al., 2022; Wang et al., 2023a; Green et al., 2012), they are trained to gather information only at the observational level, which may limit their understanding of the physical world (lack of interventional and counterfactual information). While these models, including CLIP (Radford et al., 2021) and GPT-4 (Achiam et al., 2023), excel in various tasks like vision-language understanding and image recognition with multi-modal input, they are limited by issues like hallucination (Huang et al., 2023a; Gunjal et al., 2023) and reduced multi-step reasoning abilities (Jin et al., 2024; Sun et al., 2023), indicating the lack of causal understanding. On the other hand, canonical causality research fails to incorporate multi-modal datasets and is typically tied with predefined variables for SEM. Recent work (discussed in Section 6) reveals the synergy between Causal ML and foundation models for mutual enhancement, and paves the way for FVWMs.
Misconception II: Causal ML research must be theory-oriented.
The traditional emphasis on theory-oriented approaches in causal machine learning research often overshadows empirically-driven approaches, which prioritize causal experimentation, verification, and hypothesis selection (Dorie et al., 2019; Jensen, 2019; Gong et al., 2022; Nie & Wager, 2021; Saito & Yasui, 2020; Vansteelandt et al., 2012; Schuler et al., 2018; Rolling & Yang, 2014; Matthieu & Gaël, 2023; Cook et al., 2002; Hoover, 2001; Bareinboim & Pearl, 2016; Bareinboim, 2014). Consequently, over-reliance on purely theoretical approaches can lead to models that fail to perform effectively in practical, complex scenarios. We would like to emphasize that one of the end goals of causality research is causal-aware reasoning and the ability to predict the consequences of any action accurately. Canonical formulations with theoretical guarantees are one approach to causal research, while focusing on empirical evaluations of causal properties is an alternative when it comes to FVWMs. A shift towards empirically-driven research allows models to assimilate causal knowledge from a wide array of scenarios, akin to the way large language models are trained and evaluated. Moreover, the unique problem setting of Embodied AI opens up the possibility for empirical evaluation to become the cornerstone of causal model assessment. Through interaction with both real and simulated environments, Embodied AI agents can generate and leverage interventional or counterfactual data, thereby permitting a more pragmatic approach to causal understanding. Thus, empirically-driven causal research under FVWMs and embodied AI is equally or more important than the theory-oriented paradigm.
Misconception III: Predictive models cannot be causal.
The principle that correlation does not imply causation is highly relevant in the field of predictive modeling. It is often misunderstood as meaning that predictive models cannot be causal. This interpretation holds some truth when models are trained solely on observational data, which are prone to confounding effects that can bias predictions about the outcomes of specific actions. However, it does not apply when a predictive model is trained with active intervention data, which allows the model to demonstrate causal properties. This aligns with the discussion in misconception I : RL models are inherently causal and can learn to take optimal actions for a variety of tasks (Kaddour et al., 2022; Richens & Everitt, 2024). Thus, causal models are inherently high-accuracy prediction models of consequences about all possible actions. Predictive models can therefore be considered causal models when trained with appropriate algorithms and datasets, even without explicit causal theory considerations. In the realm of Embodied AI, opportunities arise for gathering suitable interaction data in the physical world. This enables the integration of both causal considerations and large-scale prediction models (Ke et al., 2019; Scherrer et al., 2021; Toth et al., 2022).
4 Research opportunities
Recognizing the immense potential of foundation models for Embodied AI and the necessity of a causal perspective to ensure accurate estimation of action consequences, we foresee significant research opportunities, especially since current foundation models are not causal-aware. In particular, we identify a pressing need for research advancement in the following aspects.
Diverse modalities
The foundation model for Embodied AI extends beyond traditional vision and language realms. Robotics necessitates the integration and interpretation of diverse sensory data, including tactile and torque sensors, covering a broad range of modalities which vary from one agent to the next. This requirement surpasses the capabilities of current vision and language-based foundation models. Developing a system capable of effectively processing and utilizing these complex sensory data is crucial for enhancing robotic capabilities, especially in more intricate and nuanced tasks. On the other hand, building a causally-aware FWMs that comprehends high-dimensional observations from a large number of potential different modalities to model the underlying mechanisms is very challenging and presents multiple research opportunities.
First, integrating data from multiple modalities to achieve a cohesive understanding from different sensory inputs poses challenges in spatial and temporal alignment (Ward-Cherrier et al., 2018; Lin et al., 2023b; Wang et al., 2023b; Huang et al., 2023b; Behley et al., 2019) due to the varying perspectives and information overlap provided by each sensor, as well as the need to reconcile different sampling intervals to maintain an up-to-date representation of the world.
Second, agents vary widely in their sensor and effector configurations, from humanoid robots like Tesla’s Optimus to quadrupedal robots like Boston Dynamics’ Spot, to virtual cars, each with unique sensory inputs and available effectors. For FWMs to make precise action-outcome predictions, they must semantically interpret these diverse inputs and controls, necessitating a representational model that encapsulates the essence of an agent’s design to enable generalization across different configurations.
Lastly, handling the complexities of varying abstraction levels for effective interaction presents a formidable challenge (Schölkopf et al., 2021; Wang & Jordan, 2021; Koul et al., 2023; Zhang et al., 2020c; Lamb et al., 2022). In standard causal models, variables are precisely defined within structural equation models (SEMs) for a single dataset and task. However, such clarity is often absent in the realm of multi-modal, high-dimensional inputs across diverse tasks. While emerging works are exploring causal-aware foundation models (Zhang et al., 2023c; Ke et al., 2022), they remain confined to tabular data with predefined tasks. Given a high-dimensional input, a representation meaningful for interactions and their effects (affordances) is necessary to ensure the world model is meaningful. For instance, modeling interaction and consequences at a pixel level provides little immediate insight and requires additional transformations to either map the (motor) control of an agent into actions in pixel space or to extract an interpretable understanding of the effect the interaction had. Additionally, a foundational world model for embodied AI, which aims to achieve broad generalization, must consider different knowledge hierarchies to ensure effective generalization and knowledge transfer across varied contexts. For example, understanding the consequences of avoiding a car attempting to overtake should be consistent regardless of the specific car instances (car, brand), and the same principle applies to home robotics with different effectors for similar tasks. The FVWM, therefore, requires hierarchical representations that facilitate knowledge transfer at multiple levels, enabling nuanced understanding and application across various scenarios and tasks.
New paradigm for interactions
As discussed, canonical research in causality and RL shares similar objectives but differs in the source of the data: observational data versus interventional. Traditional causality research seeks to learn the world model using existing data, assuming that new interventions are not feasible, while online RL learns through interactions. For Embodied AI with FVWM, there will be access to a vast amount of offline observational data from humans, such as vision language data, human motion data, and robotics data, along with the possibility to obtain new intervention data. We need to utilize all different data paradigms, breaking down the barriers between foundation models, RL, and causality.
Pre-training FVWMs from offline data alone may be sufficient to obtain good performance in some scenarios. However, we anticipate that gathering additional online interaction data (Lee et al., 2022) will be necessary to obtain robust performance across general tasks and environments. This data should be collected in a way such that it enables the model to reason about interventional and counterfactual scenarios and therefore integrating different levels of causal reasoning (Pearl, 2009b). Therefore, a key research direction will be developing algorithms that efficiently gather the most useful intervention data online, enabling fast adaptation:
  • The related areas of active learning and experimental design selectively query datapoints to maximize information gain (González et al., 2016; Ma et al., 2019; Gong et al., 2019; Yin et al., 2020). Similar ideas have been explored in the context of training world models by collecting data where the world model has the greatest epistemic uncertainty (Sekar et al., 2020; Rigter et al., 2024). However, maximising uncertainty does not necessarily lead to the collection of data that is most useful for improving decision-making (Pathak et al., 2017). Another paradigm then is to incentivize agents to collect semantically diverse data from online interactions based on representations grounded in natural language (Gupta et al., 2022; Tam et al., 2022) or control-relevant information contained in observations (Lamb et al., 2022; Islam et al., 2022).
  • The research into unsupervised environment design (Dennis et al., 2020; Samvelyan et al., 2023) aims to learn to adapt the difficulty of the environment to the agent’s learning progress and can produce a natural curriculum of increasingly complex environments for agents to interact with and collect data from, thereby improving the knowledge of world model in a continual manner.
In summary, by training a causally aware model, FVWMs may unlock opportunities to develop algorithms that efficiently gather the intervention data that is most relevant for improving decision-making. Such an approach has the potential to greatly expand the generality and reliability of the world model while reducing the cost of online data acquisition.
Improving planning
FVWMs aim to facilitate long-term decision-making in Embodied AI systems. It is well-known that improved prediction accuracy leads to improved planning (Kearns & Singh, 2002; Janner et al., 2019). However, even with an accurate world model, optimising behaviour with large prediction models is still a challenging problem and requires either reinforcement learning (Hafner et al., 2020), search (Schrittwieser et al., 2020), or sophisticated model-predictive control (Hansen et al., 2022) within the world model. While these approaches have proven successful in narrow tasks, it is unclear whether they will scale to foundation world models trained across diverse tasks and environments. By modelling causal structure, FVWMs have the potential to learn representations that are highly suitable for planning and optimisation. Thus, investigating how the representations learnt by FVWMs can improve not only prediction accuracy and generalization, but also planning, is another key research direction.
Latent Dynamic Representations for World Models
To enable sample efficient learning and planning with strong generalization abilities and causal reasoning, it will be crucial for FVWMs to learn suitable abstract representations of the world. Current large world models such as GAIA (Hu et al., 2023), LWM (Liu et al., 2024) learn to encode the entire images into a latent space based on a reconstruction loss which can capture many irrelevant details for decision-making. Therefore, minimal state representations (Lamb et al., 2022; Islam et al., 2022), which only capture agent controllable features of the environment, are more suitable for causal reasoning and planning. Some possible approaches are using regularisation techniques such as sparsity or causal priors (Li & Han, 2024; Killedar et al., 2021) in the latent space, which can be thought of as encouraging the latent space to learn the SEM (Rajendran et al., 2024). Other possibilities include changing the model architecture to learn latent representations that disentangle task-irrelevant visual details from task-relevant details and dynamically adapt these representations as task distribution of interest changes. With improved state representations, we expect that the model should be able to capture causal relationships and therefore learn in a more sample-efficient way and generalise better.
Empirically-driven research and evaluation
The goal of Embodied AI research is to improve the performance and applicability of embodied agents in real-world scenarios. Therefore, to measure success we should focus on empirical performance on metrics related to physically meaningful tasks, rather than theoretical results. Thus, shifting the research culture towards being empirically-driven, especially for causal researchers, and developing relatable, realistic, and economical evaluation methods is crucial for success.
Evaluating real-world scenarios with a physical system is very costly in terms of time and potential risks, although it is necessary for the ultimate success of Embodied AI. Thus, constructing a curriculum of tasks that are easy to evaluate is key. This means evaluations can be performed at scale, replicated by any research institute, have clear metrics demonstrating progress, and provide a reliable approximation for end tasks in the real world. This requires not only further advancements in Embodied AI in simulated environments like CausalWorld (Ahmed et al., 2020; Bauer et al., 2022) and Habitat 3.0 (Puig et al., 2023) but also the creation of innovative tasks that aid in this process. For example, evaluation on counterfactual prediction can be a good approximation for assessing true world understanding. However, we need to consider concrete tasks and evaluation metrics. In the case of counterfactual prediction, the task could involve generating a video (Yin et al., 2023), while the metric needs to be nuanced as pixel-space evaluation has drawbacks (Fernandez-Moral et al., 2018).
Moreover, it’s unlikely that a single metric will fully capture the model’s performance, especially with the new causal perspective. Traditional metrics such as graph prediction accuracy or treatment effect estimation accuracy with well-defined variables become less relevant. However, the philosophical approach and perspective in evaluating causal properties remain crucial. For example, it’s worth considering if evaluating causal metrics in a representation space will provide additional insights complementary to prediction tasks. Also, the key to Embodied AI is understanding affordances and the consequence of actions, so evaluating counterfactuals in an imagined world might be effective.
On the other hand, the primary reason that correlation-based models are often considered unsuitable for causal-aware tasks is the bias present in the datasets collected. A causal-aware model is expected to accurately predict future outcomes in out-of-distribution evaluations (similar to A/B testing data), even when trained with biased data. Therefore, it’s crucial to consider the availability of training data and the types of properties that reflect real-world scenarios, while also ensuring that the evaluation dataset genuinely represents generalization across various scenarios.
In conclusion, while the broader concept of Embodied AI encompasses virtual and mixed reality environments, as well as physical real-world environments (robotics and human action) , the final evaluation, especially considering human safety, cannot be overlooked (Kadian et al., 2020; Ma & Zhang, 2023; Jin* et al., 2023; Yang et al., 2023). All these efficient evaluation methods aim to pave the way for real-world deployment.
5 Deploying Embodied AI
Many factors affect how FVWMs can impact the deployment of Embodied AI in the real world. In this chapter, we focus on what we believe will be one of the most important factors, which is FVWMs ability to greatly scale up the deployments of Embodied AI in robotics.
General purpose robots
General purpose robots form perhaps the most exciting prospect of generally capable Embodied AI. By compatibility with human environments and tools, robots of this kind could finally bridge the gap between the physical and digital worlds and allow large-scale automation of physical tasks. Potential applications include automating domestic tasks (Puig et al., 2023), nursing (Ohneberg et al., 2023), work in hazardous environments (Petereit et al., 2019), and manufacturing (Kheddar et al., 2019). We discuss ethical considerations regarding societal impact in our impact statement. Recent work has demonstrated zero-shot solutions to complex tasks, yet shows poor performance when faced with new environments. Research in causal FVWMs should prioritize consideration of diverse modalities and platforms, enabling success across a wide range of tasks and environments.
Rapid deployment of specialized robots
While general purpose robots capable of zero-shot learning stand to reduce deployment time, specialized robots will remain more cost effective in some scenarios - especially for large-volume production. Examples include farming (Aravind et al., 2017), scientific lab exploration (Ren et al., 2023; Boiko et al., 2023) and industrial welding machines (Grau et al., 2020). Implementing specialized robots is a time-consuming process and therefore learning to perform tasks without explicit programming an important research area (Dzedzickis et al., 2022). However, current learning implementations are typically limited to a single environment (Meyes et al., 2017). We believe that causal-aware FVWMs may resolve these constraints and therefore reduce the cost to deploy specialized industrial robots. An important testbed for this work will therefore be operation across robot morphologies with different sensing and tooling attachments.
Robustness and safety
Industrial robots have typically been confined to isolated cells. However, constraints such as force limitations have enabled more collaborative operation, improving productivity (Dzedzickis et al., 2022). In other applications, such as nursing homes, direct physical interaction with humans is vital. Yet in such situations, simple threshold constraints may not be sufficient to ensure physical safety while still allowing the robot to perform the tasks of interest. A thorough grasp of the outcomes resulting from physical interactions, as facilitated by FVWMs, presents an opportunity to enhance safety by developing more nuanced safety constraints.
6 Related Work
Foundation model for Embodied AI
There have been a number of surveys on foundation models with a focus on recent research advances, challenges, and opportunities (Zhou et al., 2023a; Bommasani et al., 2021). Regarding Embodied AI, Firoozi et al. (2023); Xiao et al. (2023) survey the role of pretrained foundation models in robotics for perception, navigation, decision making and control. Recent studies have explored the use of vision-language FMs in Embodied AI tasks to improve sample efficiency and performance abilities (Khandelwal et al., 2022; Batra et al., 2020; Shridhar et al., 2022). Moreover, numerous surveys from the perspective of robotic tasks have been published, providing insights applicable to both deep learning and machine learning at large (Yin et al., 2021; Newbury et al., 2023). Recent research advances have increasingly focused on utilizing foundation models for Embodied AI. A number of methods leverage large LLMs and VLMs for planning, re-planning and navigation (Lin et al., 2023a; Schumann et al., 2023; Skreta et al., 2024). Also, more recent work focuses on generalization across different environments and tasks for robots based on VLMs, such as SayCan (Brohan et al., 2023b), RT-2 (Brohan et al., 2023a), GATO (Reed et al., 2022), AutoRT (Ahn et al., 2024), and autonomous driving, such as GAIA (Hu et al., 2023).
Causality and Causal Foundation Model
In recent years, books (Pearl, 2009b; Hernán & Robins, 2010; Cunningham, 2021; Imbens & Rubin, 2015; Spirtes et al., 2000; Peters et al., 2017) as well as surveys (Spirtes & Zhang, 2016; Kitson et al., 2021; Shojaie & Fox, 2022; Zanga et al., 2022) have been published, drawing attention to causality and causal machine learning in general. Related to our work is the recent overview of open problems from (Kaddour et al., 2022), which connect different machine learning approaches to causality (Schölkopf, 2022); and a survey and position paper on causal reinforcement learning (Zeng et al., 2023) which aligns with our position of breaking the boundary of between these fields. There is large body of research work considering generic deep learning and causality including (Schölkopf et al., 2021; Jin et al., 2022; Zhou et al., 2023b; Berrevoets et al., 2023).
With the advent of foundation models, particularly large language models (LLMs), there have been multiple position and evaluation papers considering whether LLMs are causally aware and if they can be used to complete causal tasks (Willig et al., 2022; Tu et al., 2023; Kıcıman et al., 2023; Zhang et al., 2023a, b; Jin et al., 2023; Jin* et al., 2023; Zečević et al., 2023). Generally, these papers argue that existing LLMs can discuss causality but are not capable of reasoning in a causally aware manner. Consequently, this further confirms that current LLMs are not yet world models. Meanwhile, recent research has begun to show that it may be possible to innovate based on transformer-style architecture (Lorch et al., 2022; Zhang et al., 2023c), enabling foundation models to estimate treatment effects. Although these works are in their early stages, they provide guidance and the possibility for a world model that is causally aware, which will be essential for Embodied AI.
Causal World Models
World models, often used in RL, aim to model the transition dynamics which are expected to be intrinsically causal. It is a longstanding challenge, where early work has used standard sequence models for this purpose (Schmidhuber, 1990). More recent research has focused on using Variational Autoencoders (VAEs) to map inputs into a latent representation to enable dynamics modelling with high-dimensional observations (Ha & Schmidhuber, 2018; Hafner et al., 2019; Gelada et al., 2019; Schrittwieser et al., 2020). Advancements in the Dreamer series (Hafner et al., 2020, 2023) and transformer-based world models (Chen et al., 2022; Micheli et al., 2023; Robine et al., 2023) highlight the benefits of discrete latent variables and transformers for latent dynamics. Recent works have also leveraged diffusion models for world models (Rigter et al., 2023; Alonso et al., 2023) to improve prediction accuracy at the cost of increased computation.
The community is increasingly recognizing the important role of causal properties in world models. Recently, Richens & Everitt (2024) proved that any agent capable of generalizing under domain shifts must learn a causal model, and there have been many works trying to explicitly incorporate causality. Zhang et al. (2020b, 2019) use causal core sets (Zhang et al., 2020a) to model latent transition dynamics. Li et al. (2020); Lu et al. (2018), on the other hand, considers the confounding effect during the modelling of state transitions. Another line of work (Annabi, 2022; Volodin et al., 2020; Zhu et al., 2022; Wang et al., 2022; Tomar et al., 2021) explicitly build a SEM among variables in the states, allowing the transition dynamics to only depend on causally-relevant feature. Unfortunately, existing causal world models fall short as FWMs due to their inability to generalize across domains and tasks, and to process multi-modal inputs efficiently. Future research on causal-aware FVWM will bridge the gap and enable Embodied AI.
7 Conclusion
In this position paper, we advocate for FVWMs as key to unlocking the full potential of embodied AI in both general and scalable contexts, highlighting the indispensability of causality in these models. We reiterate that causality as a property is fundamental to accurate action and consequence prediction, and its integration is crucial in the realm of machine learning and foundation models. Looking forward, we envision future research focusing not only on high-modality observations but also on transcending the boundaries of various machine learning sub-fields. This approach aims to optimize data utilization through both physical interactions and a blend of online and offline methodologies. Moreover, we emphasize the significance of establishing empirical evaluation benchmarks to foster real-world impact in a cost-effective and low-risk manner.
While this paper concentrates on the FVWMs in embodied AI, we acknowledge that numerous other research challenges remain unaddressed by this study. These include, but are not limited to, advancements in hardware, human-robot communication and collaboration, and comprehensive socio-economic considerations, which are very important for embodied AI in general.
Impact statement
This position paper, focusing on world models for Embodied AI, does not propose new technological advancements with direct societal consequences in isolation. However, it is crucial to address the broader social implications of the field. This is a topic of intense debate with a multitude of works predicting its outcome and proposing how to mitigate risks. However, a reoccurring theme in the debate is that is will be necessary to ensure an equitable transition of society to rapidly increasing automation. Willcocks (2020) argues that such a transition will unlikely result in a net job loss, and will be vital to mitigate changing demographics which are predicted to lead to labour shortages. Using foundation models in Embodied AI shifts the focus of automation from tasks that are dirty, dull, and dangerous (3D) to those that necessitate advanced cognitive and reasoning skills. As a result, the implications of widespread deployment in automating various job types could vary significantly, necessitating careful consideration by society.
Another important consideration is that Embodied AI should be ethical and fair. For the real-world evaluation of embodied AI, safety and ethical considerations are key, as there is the possibility of inducing physical harm. Moreover, Causal FVWMs may reduce biases and lead to fairer algorithms (Kusner et al., 2017; Zhang & Bareinboim, 2018). However, as suggested by Nilforoshan et al. (2022), this may not directly translate to equitable decision making. Instead it will remain important for researchers to develop this area further.
Acknowledgment
The authors wish to express gratitude to Katja Hofmann, Jan Peters, Nan Duan, Anastasia Varava, Xin Tong, Tom Minka, Jiang Bian, David Sweeney Baining Guo and John Langford, whose discussions have inspired and aided the manuscript.